Find an equation of the line that satisfies the given condition. (Let x be the independent variable and y be the dependent variable.)
The line passing through the origin and parallel to the line joining the points
(4, 5) and (6, 6)

Answer by TimothyLamb(4379)

---
given:
line through poitns: (4, 5) and (6, 6)
---
slope of parallel line:
m = (6-5)/(6-4)
m = 0.5
---
given:
line through the origin: (0,0)
---
to solve the linear equation, copy and paste this (the given point and parallel slope):
0,0,0.5
into the "Point-Slope form: x1 y1 m" input box here: https://sooeet.com/math/linear-equation-solver.php
---
x-intercept = (0, 0)
y-intercept = (0, 0)
slope = 1/2 = 0.5
---
answer:
the slope-intercept form of our new parallel line:
y = (1/2)x
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php